OpenStudy (anonymous):

Find the missing coordinate for the given slope and points. m = -2 (3, 5), (4, __)

OpenStudy (bibby):

y2-y1 m=----- x2-x1

OpenStudy (nirmalnema):

y2-y1 / x2-x1 = m (5-y1) / (4-3) = -2 5-y1=-2 so y1 = 7